# Probing the fission of 220Ra⋆ at E⋆ 31.8–45.4 MeV

### v1

@@ Abstract
In this work, fission of moderately excited compound nucleus 220Ra⋆ produced in 12C+208Pb reaction at Elab = 81.9, 75.8, and 67.5 MeV has been studied. 28 fission fragments, within the mass range 60A141, have been identified based on their characteristic decay -lines and half-lives. The yields of different fission fragments have been analyzed to generate isotopic and isobaric yield distributions. The value of the mass dispersion parameter, , is found to be 2.93 and 2.65 for Antimony (Sb) isotope at excitation energy E⋆ = 45.4 and 39.6 MeV, and 1.24 for Indium (In) isotope at E⋆ = 45.4 MeV. The charge dispersion parameter for Sb is estimated to be 0.769 and 0.714 at E⋆ = 45.4 and 39.6 MeV, respectively. For In isotopes, the value of is estimated to be 0.430 at E⋆ = 45.4 MeV. The value of mass and charge dispersion parameters for Sb and In isotopes display good agreement with the values reported in the literature for similar systems. It has been found that 99mTc and 111In, medically important isotopes, populate in this system. The present investigations suggest that fission is a dominating mode of deexcitation of the compound nucleus, even at entrance channel energies slightly above the Coulomb barrier.
- September 22, 2024
@@ 1 Introduction
Nuclear fission dynamics has been the subject of rigorous investigations over the past few decades because of the interdisciplinary outcome potential of the process [1]. The cross-section data generated for fission fragments in reactions with various projectile and target combinations are inputs for research in super-heavy element synthesis [2, 3], nuclear astrophysics, medical applications, and security purposes. The cross-section data is vital for developing next-generation nuclear reactors for producing nuclear energy, producing new radioactive isotopes, and transmuting commercial nuclear waste [4], to name a few. The mechanism of fission manifests only when the fissioning nucleus surmounts the fission barrier [5], which emerges from the interplay between Coulomb and surface energy factors in the semi-empirical formula. No fission barrier exists for Z 104; therefore, a symmetric, single-humped distribution of masses of fission fragments is most probable. However, the observation of asymmetrical mass distribution in fission involving actinide targets [6] and synthesis of elements beyond Z=104 [7] hinted at the discrepancy in the liquid drop model. The asymmetry in the mass distribution was later explained based on nuclear shell effects by incorporating the shell correction term in the liquid drop model energy [8, 9, 10].
@@
Additionally, the increased excitation energy decreases the mass distribution asymmetry of the fission fragments, leading to symmetric fission. This feature may be explained as a result of a gradual decrease of shell effects with an increase in excitation energies of the compound nucleus as reported in [11, 12]. Although a combination of the shell and liquid drop models offers insight into the fundamental aspects of nuclear fission, certain aspects still need to be understood. As a result, nuclear fission continues to be an important area of investigation in nuclear physics. The effect of excitation energy and angular momentum on various fission observables has provided more profound insights into the mechanism of fission. In Refs. [13, 14], the mass-asymmetry, shape-deformation, and coulomb factor (ZPZT) have been found to affect the dynamics of fusion-fission. The entrance-channel spectroscopic properties of 12C+208Pb system, along with the similar systems studied elsewhere, are presented in Table 1 for ready reference.
Hinde . [15] studied the angular distributions of fission fragments for 16O+238U system. Ghosh . [16] measured the mass distribution of fission fragments for deformed and spherical targets with light projectiles, considering the microscopic effects arising from the orientation of the deformed targets to obtain insights into the different paths followed by the excited composite system towards fission. It has been found that, for a deformed target, the width of the mass distribution of fission fragments from fusion-fission reactions sharply increases with increased excitation energy around the Coulomb barrier. In contrast, there is a smooth variation in the width of the mass distribution for a spherical target at the same energies.
Experimental studies have consistently shown that actinides with masses up to A 256 mainly undergo asymmetric fission [17], whereas nuclei around 208Pb offer symmetric fission [18]. In a study performed by Itkis . [19], for the fission of Ra isotopes produced in 12C+204,206,208Pb reactions, the effect of compound nucleus excitation energy on the ratio of symmetric to asymmetric component yield has been observed. In Ref. [20], three fission modes of compound nuclei 220Ra formed in 12C+208Pb reaction were experimentally observed. Schmidt . [21] studied 70 systems from 205At to 234U via Coulomb-excitation induced fission. It has been found that the transition takes place from the symmetric to asymmetric fission mode around mass A226 in this region of nuclei. In Ref. [22] the entrance channel effect in the fission of 216Ra has been studied from the mass and energy distributions of fission fragments from two reactions 12C + 204Pb and 48Ca + 168Er, the contribution of asymmetric fission was only around 1.5% but for the latter was around 30%. The observation of unexpected asymmetric fission of 180Hg (after decay) [23], sparked a profound curiosity in theoretical modeling [24, 25, 26, 27]. Recent experimental work [28, 29] shows that in the sublead region, asymmetric mode is present at moderately high excitation energies. Karapetyan . [30] studied the interaction of 11B with 209Bi and 181Ta targets. In this study, the presence of the asymmetric component for 209Bi at 146 MeV was observed, whereas it is absent for 181Ta at 125.7 and 245.4 MeV. Miernik [31] reported the fission of 215Fr, the presence of an asymmetric component in the charge distribution at an excitation energy of 61.0 MeV. Collectively, there is no unambiguous explanation for the dependence of fission on excitation energy, whether the microscopic structure is preserved, and to what excitation energies.
@@
To supplement the ongoing efforts toward a comprehensive understanding of fission dynamics, an experiment has been carried out to measure the production cross-sections of fission fragments produced in the 12C+208Pb reaction at Elab = 81.9, 75.8, and 67.5 MeV. The experimental methodology is outlined in Section 2, while Section 3 presents the details of data analysis. Section 4 delves into the results and discussion. Lastly, Section 5 summarizes the results and conclusions.
2 Experimental Mehodology
@@
The experiment was carried out at the Inter-University Accelerator Centre (IUAC), New Delhi, India, in the General Purpose Scattering Chamber (GPSC), equipped with an in-vacuum transfer facility (ITF) for swift transfer of irradiated samples from the chamber to the detection setup. Enriched 208Pb (99.74%) was utilized to prepare the target foils of thickness 0.169-0.237 mg/cm2. 208Pb were deposited on Al foils (thickness range of 1.0–1.5 mg/cm2) using high vacuum evaporation technique [36]. The Al foils were prepared using the cold rolling machine, and their thickness was measured using the transmission method. Fig. 1 depicts a stack containing 208Pb target and Al-catcher foils (3.0–4.5 mg/cm2), mounted on an aluminum target ladder with a concentric hole of 1.2 cm diameter in GPSC for irradiations. Al-catcher foils were positioned in front and behind the target foil to trap fission fragments recoiled out of the target in forward and backward cones.
@@
The irradiations were performed using 12C beams obtained from 15-UD pelletron accelerator at energies Elab (corrected) = 81.9, 75.8, and 67.5 MeV with the beam current ranging between 9–10 nA for 6–9 hours depending upon the half-lives of residues. The beam current was monitored using an ORTEC current integrator installed downstream of the beamline behind the target-catcher assembly. The energy-loss corrected beam energy, i.e., the effective incident beam energy at the half-thickness of the target in each target-catcher foil assembly, was estimated using the code SRIM [37].
@@
After the irradiation, the target-catcher assembly was taken out of the scattering chamber using an ITF and moved to the offline counting setup. The -ray activities induced in the target-catcher assembly were counted with a pre-calibrated HPGe clover detector coupled to a CAMAC-based DAQ (Data Acquisition) system running on locally developed CANDLE software [38]. The schematic representation of this setup is shown in Fig. 2. For calibration purposes and to estimate the geometry-dependent efficiency of the detectors, standard sources, i.e., 152Eu, 60Co, and 133Ba, with known strengths were used. The measurements with the standard sources were taken before, in between, and after counting the irradiated target-catcher assembly at different source-detector distances to monitor the change in the efficiency of the system during its operating period by Clover detector, if any. The resolution of the clover detectors was estimated to be 2.3 keV at the 1332 keV -ray of the 60Co source. The target-catcher assemblies were counted in the same geometry as the standard -ray sources to avoid the solid angle effect during the counting. During the counting, the dead time of the detector was kept 10%.
| S. No. | Eγ (keV) | Iγ() | t1/2 | Nuclide |
| 1 | 1332.5 | 88 | 23.7 min | 60Cu |
| 2 | 596.56 | 26 | 9.19 h | 62Zn |
| 3 | 381.85 | 28.3 | 2.26 h | 66Ge |
| 108.85 | 10.6 | 2.26 h | 66Ge | |
| 4 | 167.01 | 84 | 18.9 min | 67Ge |
| 5 | 668.21 | 22.1 | 52.6 min | 70As |
| 6 | 834.13 | 95.45 | 14.10 h | 72Ga |
| 7 | 406.5 | 12.1 | 14.8 h | 76Kr |
| 8 | 589 | 39 | 106.3 min | 80Sr |
| 9 | 402.58 | 50 | 76.3 min | 87Kr |
| 10 | 140.51 | 89 | 6.01 h | 99mTc |
| 11 | 306.83 | 89 | 14.2 min | 101Tc |
| 12 | 469.34 | 18.31 | 4.44 h | 105Ru |
| 724.21 | 47.8 | 4.44 h | 105Ru | |
| 13 | 302.77 | 66 | 21.7 min | 107Rh |
| 14 | 203.3 | 74.2 | 4.16 h | 109In |
| 15 | 657.75 | 98 | 4.92 h | 110In |
| 16 | 245.35 | 94.1 | 2.80 d | 111In |
| 17 | 497.31 | 97.9 | 32.1 min | 115Sb |
| 18 | 93.7 | 33.1 | 2.49 h | 116Te |
| 19 | 1293.55 | 100 | 60.3 min | 116mSb |
| 20 | 158.56 | 85.9 | 2.80 h | 117Sb |
| 21 | 273.34 | 28 | 2.49 h | 117Cd |
| 22 | 719.7 | 64.7 | 62 min | 117Te |
| 23 | 1229.65 | 100 | 5.00 h | 118mSb |
| 24 | 536.06 | 99 | 12.36 h | 130I |
| 25 | 529.87 | 87 | 20.83 h | 133I |
| 26 | 847.02 | 96 | 52.5 min | 134I |
| 27 | 249.79 | 90 | 9.14 h | 135Xe |
| 28 | 190.32 | 44.8 | 18.27 min | 141Ba |
@@ 3 Data Analysis
The fission residues were identified using the characteristic -rays and their half-lives obtained from the decay curve analysis. The residues of longer half-lives were counted for a week or more. The -ray spectrum of 12C+208Pb system obtained at Elab = 81.9 MeV is presented in Fig. 3 in which the -lines associated with different fission fragments are marked. The inset shows the decay curve of 141Ba residue obtained following the 190.32 keV -line. As shown in this figure, the activity reduces to half of its value in 18.27 min of lapse time, which shows good agreement with the known characteristic half-live of 141Ba and confirms its identification. The same procedure has been followed for all the fission fragments identified in this work at different energies. In Table 2, the fission fragments with their spectroscopic properties, e.g., the half-lives, -ray energies, and intensities are taken from the Decay Radiation database within the NUDAT (Nuclear Structure and Decay Data) application [39] and ref.[40], are given. Moreover, it should be mentioned that the activation technique used in the present work imposes some restrictions on identifying reaction residues due to its limitations. For example, it is not possible to measure stable, very short-lived, or very long-lived isotopes in such experiments.
| Cross-sections of fission fragments, (mb) | ||||
| Nuclide | 45.4 MeV | 39.6 MeV | 31.8 MeV | |
| 60Cu | 44.734.71 | 30.494.12 | ||
| 62Zn | 88.717.98 | 41.384.33 | 33.823.46 | |
| 66Ge | 63.455.47 | 37.413.78 | 28.472.86 | |
| 67Ge | 60.385.59 | 34.703.79 | ||
| 70As | 69.306.96 | 41.735.46 | 33.72 | |
| 72Ga | 73.866.16 | 39.613.81 | ||
| 76Kr | 63.036.21 | |||
| 80Sr | 48.064.43 | 23.242.56 | ||
| 87Kr | 14.211.43 | 8.291.19 | ||
| 99mTc | 12.541.19 | 7.910.77 | 5.600.48 | |
| 101Tc | 48.344.79 | |||
| 105Ru | 49.564.45 | 32.713.25 | 24.162.40 | |
| 107Rh | 46.204.50 | |||
| 109In | 31.792.75 | |||
| 110In | 102.4910.47 | 51.484.69 | 36.343.27 | |
| 111In | 50.824.48 | |||
| 115Sb | 36.953.52 | 20.16 | ||
| 116Te | 68.265.86 | 33.863.16 | 16.441.62 | |
| 116mSb | 72.496.66 | 43.285.17 | ||
| 117Sb | 39.723.50 | 22.091.97 | 12.221.13 | |
| 117Cd | 63.465.62 | 17.481.93 | ||
| 117Te | 77.096.96 | |||
| 118mSb | 5.220.62 | 4.240.52 | ||
| 130I | 17.761.64 | |||
| 133I | 20.962.00 | |||
| 134I | 42.544.01 | 29.293.18 | 19.232.06 | |
| 135Xe | 39.004.18 | |||
| 141Ba | 31.483.29 | 20.972.82 | ||
@@
After identification of all fission fragments, their production cross-sections () were calculated using the standard activation equation given elsewhere [41]. It may be pointed out that the value of measured in the present work for respective residues is cumulative because these cross-sections include the contribution from the decay of the neighboring unstable isobars. The error in the production cross-sections may arise due to the uncertainty in target thickness, beam flux, and the geometry-dependent efficiency of the detector. The overall error in the measured cross-sections, including the statistical errors, is estimated to be . In the present work, 28, 19, and 9 fission fragments have been identified at E⋆ = 45.4 MeV (Elab = 81.9 MeV), E⋆ = 39.6 MeV (Elab = 75.8 MeV) and E⋆ = 31.8 MeV (Elab = 67.5 MeV), respectively. The production cross-sections of fission fragments, along with the errors, are given in Table 3.
4 Results and Discussion
4.1 Mass distribution of fission fragments
The mass distribution of fission fragments is a crucial post-fission observable directly linked to the collective dynamics of the fission as it describes how nuclear shell and pairing effects evolve with change in excitation energy. To generate mass distributions, the experimental cross-sections of fission fragments produced in the 12C+208Pb system are plotted as a function of mass number (A) in Figs. 4(a)- 4(c) at different energies. As can be seen from this figure, the fission fragments exhibit a broad mass distribution, ranging from A = 60 to 141, with production probabilities ranging from 4 - 102 mb, indicating the production of the fission fragments from the decay of excited compound nucleus in a fusion-fission reaction. The mass distribution of fission fragments in 220Ra contains two groups of residues, one light mass fragment group clustered around A 70 and another heavy mass fragment group around A 116, which is a mixture of symmetric and asymmetric fission. The presence of asymmetric fission suggests the role of nuclear shell effects at the studied energies. In ref.[30], it has been reported that the mass-yield distribution of 220Ra consists of one symmetric and two asymmetric components at excitation energy E* = 118.83 MeV. In a comprehensive study by Schmidt . [21], the phenomenon of fission has been investigated by measuring the fission fragment nuclear charge distributions, e.g., 209-219Ra. In this study, it has been concluded that the transition from symmetric to asymmetric fission occurs around 214Ra, 215-219Ra isotopes display measurable asymmetric components. It may not be out of order to mention that the mass distribution presented in this work shows the similarity with neighboring Ra isotopes. However, the mass distribution of fission fragments is not exhaustive due to the limitation of the experimental technique used in the present work.
In nuclear reactions where the compound nucleus is formed with excitation energy (E*) well above the fission barrier, the input angular momentum () strongly influences the mass distribution of the fission fragments. As the angular momentum increases, the saddle point becomes compact, lowering the fission barrier. This reduction in fission barrier height increases the fission probability at a particular excitation energy. The fission dynamics are significantly impacted by the predominant direction of mass flow within the dinuclear system. As reported in the literature [42, 43, 44], if the mass asymmetry () is greater than the Businaro-Gallone critical mass asymmetry (BG) [45], mass flows from the projectile to the target, leading to the formation of the compound nucleus which may eventually decay via the process of fission. If the value of , the mass flow occurs from the target to the projectile, and a dinuclear system is formed, decaying before equilibrating in all degrees of freedom. For 12C + 208Pb system, the critical mass asymmetry () is 0.86. In contrast, the mass asymmetry () is 0.89, i.e., , resulting in the occurrence of fission via compound nucleus formation path.
Further, to understand how the probabilities of the individual fission fragments change with the incident projectile energies, the cross-sections of fission fragments common at Elab = 81.9, 75.8, and 67.5 MeV energies have been plotted as a function of incident energies. Fig. 5 shows the excitation function of 9 fission fragments, i.e., 62Zn, 66Ge, 70As, 99mTc, 105Ru 110In, 116Te, 117Sb and 134I produced in 12C+208Pb system. As seen from this figure, the cross-sections of all fission fragments increase with energy, indicating the higher production probabilities of these residues at higher excitation energies, as expected.

### v3

# Investigating the fission-like fragments in the 12C + 208Pb system at E⋆ 31.8–45.4 MeV

@@ Abstract
In this work, the cross-sections of 25 fission-like fragments within the mass range 76A141, expected to be populated via fission of moderately excited compound nucleus produced as a result of complete and/or incomplete fusion in 12C+208Pb reaction at Elab = 81.9 and 75.8 MeV, have been measured using activation technique followed by offline -ray spectroscopy. The yields of different fission-like fragments have been analyzed to generate isotopic and isobaric yield distributions. The value of the mass dispersion parameter, , is found to be 2.93 and 2.65 for Antimony (Sb) isotope at excitation energy E⋆ = 45.4 and 39.6 MeV, and 1.24 for Indium (In) isotope at E⋆ = 45.4 MeV. The charge dispersion parameter for Sb is estimated to be 0.769 and 0.714 at E⋆ = 45.4 and 39.6 MeV, respectively. For In isotopes, the value of is estimated to be 0.430 at E⋆ = 45.4 MeV. The value of mass and charge dispersion parameters for Sb and In isotopes have been found to be in good agreement with the values reported in the literature for similar systems. The mass distribution of fission-like fragments is found to be fitted with a Gaussian function, except for a few data points, indicating their population via compound nucleus fission. Further, the mass variance () displays linear increment with an increase in excitation energy. Two medically important isotopes, 99mTc and 111In, are populated in this system, suggesting a potential formation route.
- December 14, 2024
@@ 1 Introduction
Nuclear fission in heavy-ion (HI) induced reactions has been extensively investigated, generating substantial experimental data across various fission systems. A better understanding of fission dynamics offers crucial insights into various fundamental nuclear processes [1]. In general, the nuclear reactions involving heavy-mass targets result in the formation of a fully equilibrated compound nucleus (CN), which may decay via light-nuclear particles and/or fission [2, 3, 4, 5]. In fission, charge and mass distributions are two crucial experimental observables that are closely linked to the collective dynamics of fission and serve as a valuable testing ground for different theoretical models. An important finding from early investigations was the presence of asymmetric mass distributions in the low-energy fission of most actinides. The asymmetry in the mass distribution was later explained based on nuclear shell effects by incorporating the shell correction term in the liquid drop model [6, 7]. Exploring the relationship between excitation energy (E*) and fission fragment mass distribution (FFMD) is crucial, as variations in energy strongly influence the resulting mass distribution. The increased excitation energy decreases the mass distribution asymmetry of the fission fragments, leading to symmetric fission. This feature may be explained as a result of a gradual decrease of shell effects with an increase in excitation energies of the compound nucleus as reported in [8, 9]. Although a combination of the shell and liquid drop model offers insight into the fundamental aspects of nuclear fission, certain aspects still need to be understood. As a result, nuclear fission continues to be an important area of investigation in nuclear physics [10, 11, 12]. Extensive research on charge, mass, energy, and angular distribution of fission fragments reported in literature [13, 14, 15, 16] has provided more profound insights into the mechanism of fission. Dispersion parameters derived from FFMD provide insights into how the mass distribution varies with changing excitation energy [17, 18]. Dispersion parameters have been obtained for various target-projectile combinations such as 19F + 169Tm [19], 11B + 232Th [20], 20Ne + 208Pb [21], 19F + 175Lu [22], 12C + 181Ta [23], 12C + 197Au [24], 11B + 209Bi and 181Ta [25], 20Ne + 232Th [26] and 19F + 232Th [27], and their mass variance behavior as a function of excitation energy, mass-asymmetry and angular momentum have also been reported in literature. It is evident from the literature that extensive research has been conducted on fission fragment (FF) dynamics within the actinide region, but the studies on fission fragment (FF) dynamics in the preactinide region are limited. Therefore, we have aimed to contribute valuable insights into the reaction mechanism involving 12C + 208Pb system at energies Elab = 81.9 and 75.8 MeV. The entrance-channel properties of 12C + 208Pb system, along with the similar systems studied elsewhere, are presented in Table 1 for ready reference.
@@
It is important to mention here that Pokrovsky et al. [31] reported the fission fragment mass yields (Y), as well as the dependence of the fission fragment total kinetic energy (TKE) and its dispersion on the fission fragment masses of the 220Ra compound nuclei produced in the 12C + 208Pb at Elab = 57, 59 and 90 MeV. The velocities and coordinates of pair fission fragments were measured using a two-arm time-of-flight spectrometer CORSET. It has been reported that at the two lowest energies, the fission fragment mass distributions are primarily symmetric and can be represented by a Gaussian function for masses between 110 and 125, along with their complementary masses. However, an asymmetric fission mode appears as ”shoulders” in the distribution. At an energy of Elab = 90 MeV, the properties of the mass-energy distributions (MEDs) resemble the liquid drop model (LDM) predictions at high temperatures, where Y(M) is Gaussian and TKE(M) forms a parabolic curve. In the present work, the production cross-sections of residues likely to be populated through fission of CN formed via complete fusion and/or incomplete fusion on the 12C + 208Pb system have been measured employing the recoil catcher activation technique followed by offline -ray spectroscopy. Data analysis has been done to obtain isotopic yield distribution and mass distribution of fission-like fragments.
The present paper is organized as follows: Section 2 describes the experimental methodology, Section 3 provides details on data analysis and interpretation of results, and Section 4 presents the summary and conclusions.
2 Experimental Methodology
@@
The experiment was carried out at the Inter-University Accelerator Centre (IUAC), New Delhi, India, in the General Purpose Scattering Chamber (GPSC), equipped with an in-vacuum target transfer facility for swift transfer of irradiated samples from the chamber to the detection setup. Enriched 208Pb (99.74%) was utilized to prepare the target foils of thickness 0.169–0.237 mg/cm2. 208Pb were deposited on Al foils (thickness range of 1.0–1.5 mg/cm2) using high vacuum evaporation technique [32]. The Al foils were prepared using the cold rolling machine, and their thickness was measured using the transmission method. Fig. 1 depicts a stack containing 208Pb target and Al-catcher foils (3.0–4.5 mg/cm2), mounted on an aluminum target ladder with a concentric hole of 1.2 cm diameter in GPSC for irradiations. Al-catcher foils were positioned in front and behind the target foil to trap reaction products recoiled out of the target in forward and backward cones.
@@
The irradiations were performed using 12C beams obtained from 15-UD pelletron accelerator at energies Elab (corrected) = 81.9 and 75.8 MeV with the beam current ranging between 9–10 nA for 6–9 hours depending upon the half-lives of residues. The beam flux was monitored with a current integrator by measuring the charge collected in the Faraday cup positioned behind the target-catcher assembly. The energy-loss corrected beam energy, i.e., the effective incident beam energy at the half-thickness of the target in each target-catcher foil assembly, was estimated using the code SRIM [33].
@@
After the irradiation, the target-catcher assembly was taken out of the scattering chamber and moved to the offline counting setup. The -ray activities induced in the target-catcher assembly were counted with a pre-calibrated HPGe clover detector coupled to a CAMAC-based DAQ (Data Acquisition) system running on locally developed CANDLE software [34]. The schematic representation of this setup is shown in Fig. 2. For calibration purposes and to estimate the geometry-dependent efficiency of the detectors, standard sources, i.e., 152Eu, 60Co, and 133Ba, with known strengths were used. The measurements with the standard sources were taken before, in between, and after counting the irradiated target-catcher assembly at different source-detector distances to evaluate detector efficiency. The resolution of the clover detectors was estimated to be 2.3 keV at the 1332 keV -ray of the 60Co source. The target-catcher assemblies were counted in the same geometry as the standard -ray sources to avoid the solid angle effect during the counting.
@@ 3 Data Analysis
The fission-like fragments were identified using the characteristic -rays and their half-lives obtained from the decay curve analysis. The residues of longer half-lives were counted for a week or more. The -ray spectrum of 12C + 208Pb system obtained at Elab = 81.9 MeV is presented in Fig. 3 in which the -lines associated with different fission-like fragments are marked. The inset shows the decay curve of 141Ba residue obtained following the 190.32 keV -line. As shown in this figure, the activity reduces to half of its value in 18.27 min of lapse time, which shows good agreement with the known characteristic half-live of 141Ba and confirms its identification. The same procedure has been followed for all the fission-like fragments identified in this work at different energies. In Table 2, the fission-like fragments with their spectroscopic properties, e.g., the half-lives, -ray energies, and intensities taken from the Decay Radiation database within the NuDat (Nuclear Structure and Decay Data) application [35] are given.
@@
After identification of all fission-like fragments, their production cross-sections () were calculated using the standard activation equation given elsewhere [36]. The error in the production cross-sections may arise due to the uncertainty in target thickness, beam flux, and the geometry-dependent efficiency of the detector. The overall error in the measured cross-sections, including the statistical errors, is estimated to be . In the present work, 25 and 19 fission-like fragments have been identified at E⋆ = 45.4 MeV (Elab = 81.9 MeV) and E⋆ = 39.6 MeV (Elab = 75.8 MeV) respectively. The production cross-sections of fission products, along with the errors, are given in Table 2. Further, an attempt has been made to study the isotopic yield distribution and mass distribution of fission-like fragments in the following sub-sections.
| S. No. | Eγ (keV) | Iγ() | t1/2 | Nuclide | (mb) | |
| 45.4 MeV | 39.6 MeV | |||||
| 1 | 406.5 | 12.1 | 14.8 h | 76Kr | 10.521.02 | |
| 2 | 589.0 | 39 | 106.3 min | 80Sr | 13.871.27 | 11.651.28 |
| 3 | 153.54 | 34 | 22.3 min | 81Sr | 14.501.32 | 12.001.26 |
| 4 | 402.58 | 50 | 76.3 min | 87Kr | 16.801.65 | 14.401.19 |
| 5 | 743.36 | 93.1 | 16.75 h | 97Zr | 32.583.37 | 23.072.56 |
| 6 | 140.51 | 89 | 6.01 h | 99mTc | 12.541.19 | 7.910.77 |
| 7 | 306.83 | 89 | 14.2 min | 101Tc | 48.344.72 | 33.763.31 |
| 8 | 724.21 | 47.8 | 4.44 h | 105Ru | 49.564.79 | 32.713.25 |
| 9 | 302.77 | 66 | 21.7 min | 107Rh | 46.204.50 | |
| 10 | 203.3 | 74.2 | 4.16 h | 109In | 31.792.75 | |
| 11 | 657.75 | 98 | 4.92 h | 110In | 102.4910.47 | 65.226.46 |
| 12 | 245.35 | 94.1 | 2.80 d | 111In | 50.824.48 | 37.273.52 |
| 13 | 497.31 | 97.9 | 32.1 min | 115Sb | 36.953.52 | 20.162.78 |
| 14 | 931.84 | 24.8 | 15.8 min | 116Sb | 72.496.66 | 43.285.17 |
| 15 | 158.56 | 85.9 | 2.80 h | 117Sb | 39.723.50 | 22.091.97 |
| 16 | 273.34 | 27.9 | 2.49 h | 117Cd | 55.194.88 | |
| 17 | 1050.69 | 97 | 5.00 h | 118mSb | 5.220.62 | 4.240.52 |
| 18 | 751.7 | 3.1 | 11.00 min | 124Ba | 44.544.07 | 32.503.15 |
| 19 | 1114.3 | 38 | 2.10 h | 127Sn | 37.393.90 | 27.742.74 |
| 20 | 536.06 | 99 | 12.36 h | 130I | 31.992.94 | |
| 21 | 529.87 | 87 | 20.83 h | 133I | 30.292.88 | |
| 22 | 847.02 | 96 | 52.5 min | 134I | 42.544.01 | 27.622.99 |
| 23 | 249.79 | 90 | 9.14 h | 135Xe | 28.093.01 | 14.331.54 |
| 24 | 108.90 | 32 | 50.65 min | 136Nd | 25.642.41 | 16.311.96 |
| 25 | 190.32 | 44.8 | 18.27 min | 141Ba | 15.041.57 | 12.981.68 |